i Review | Constants | Periodic Table Calculate the percent ionization of hydrazoic acid (HN3) in solutions of each of the foPart B 0.100 M Express your answer using two significant figures. od 0 0 ? ΑΣΦ

Part C 4.18x10-2 M. Express your answer using two significant figures. IVO A o 2 ? ΑΣΦ

Ka= 1.9*10^-5
Answer #1

- N3 + H₂O A) N3h + H2O 0.424 M (initial) 0.424 a 0.424 d 0.424 (1-0) (at equilibrium) Ka = 10.4248) (0.424x) = 0.42442 0.424B) Similarly: Ka=0.10002 1.9x105 = 0.10002 az IAX Low = 19 x 10? 0. leo. a = (1.981004)% = 1.38*10-2 % ionization = xx 100% =